# Partial Search in a Frozen Network is Enough to Find a Strong Lottery Ticket

### v1

@@ 3.3 Optimal Positioning of the Search Space
Here we perform a preliminary investigation of the optimal positioning for each given desired SLT sparsity by exploring different pruning:locking proportions with a fixed freezing ratio of the network. Figure 4 explores different configurations of a $80\%$ freezing ratio—a search space size of $20\%$ of the original dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10. As expected, the best-performing SLT of each prune:lock configuration is found at the center of the search space, which maximizes the number of candidate subnetworks. In other words, within a search space size $S$, the optimal SLT sparsity is $k=P_{r}+S/2$, where $P_{r}=\frac{\sum_{l}p^{(l)}_{p}n^{(l)}}{\sum_{k}n^{(k)}}$ is the global pruning ratio of the network. Conversely, for a given freezing ratio $F_{r}$ of the network and a desired SLT sparsity $k$, the optimal search space position is set by the pruning ratio $P_{r}=k-(1-F_{r})/2$ and the corresponding locking ratio $L_{r}=F_{r}-P_{r}$. In the cases where this would result in $P_{r}\!<\!0$ or $L_{r}\!<\!0$, we choose a best-effort approach that keeps the desired freezing ratio and sets the search space bounds to $P_{r}\!=\!0$ or $L_{r}\!=\!0$, respectively.
@@
Figure 6: Freezing achieves better accuracy to search space trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks. SLT sparsity set according to highest accuracy in Figure 5 with dense source networks.
@@ 4 Experiments
This section demonstrates that partial SLT search reduces the search space in a broad range of situations by evaluating it on image classification and node classification. We evaluate the impact of the freezing ratio on various network architectures, identifying three scenarios. Then, we explore trade-offs between accuracy, search space size, and model memory size for different network widths.
@@ 4.1 Experimental Settings
The proposed partial SLT search is evaluated on image classification using the CIFAR-1010{10}10 (Krizhevsky, 2009) dataset, and on node classification using the OGBN-Arxiv (Hu et al., 2020) dataset. CIFAR-1010{10}10 train data is split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split. We test the models with the best validation accuracy, and report the mean of three experiment repetitions.
For image classification we employ the VGG-like Conv6 (Simonyan & Zisserman, 2014; Ramanujan et al., 2020) and ResNet-1818{18}18 (He et al., 2016) architectures, and for graph node classification the GIN (Xu et al., 2019) architecture constructed with 4 layers modified by Huang et al. (2022), all implemented with no learned biases. ResNet-1818{18}18 and GIN use non-affine Batch Normalization (Ioffe & Szegedy, 2015), while Conv6 has no normalization layers. Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution (He et al., 2015).
SLTs are searched using an extension of Edge-Popup (Ramanujan et al., 2020) that enforces the desired sparsity globally instead of per-layer (Sreenivasan et al., 2022). On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent (SGD) with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively. On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW (Loshchilov & Hutter, 2019) with weight decay $0.0001$ and initial learning rate $0.01$. All experiments use cosine learning rate decay (Loshchilov & Hutter, 2017).
Figure 7: Freezing achieves better accuracy to memory size trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks.
@@ 4.2 Reducing the Search Space with Freezing
Applying the much larger ResNet-1818{18}18 to the same task results in much stronger overparametrization. Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting of much higher pruning than locking. Even in this scenario, $90\%$ sparse SLTs found in frozen networks achieve similar accuracy ($<1$ point difference) while reducing the search space by as much as $80\%$.
As an example of the scenario where optimal SLT sparsities are found in the denser range, benefitting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin. Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the search space our method finds better performing SLTs of $70.2\%$ accuracy with $40\%$ sparsity.
These results suggest that parameter freezing at an appropriate ratio has the effect of eliminating local optimal solutions from the search space. It is known that Edge-Popup struggles to find the SLTs of the global optimal solution (Fischer & Burkholz, 2022). While searching for SLTs in a dense network leads to convergence to a local optimal solution, a moderate reduction of the search space may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space. In other words, we conjecture that if the network is properly frozen, the local optimal solution for the reduced search space is close to the global optimal solution of the entire search space.
4.3 Accuracy to Search Space Trade-Off
@@
Compared to the dense source network, the results show in all cases that, when the frozen network is twice as wide, the found SLTs for a given sparsity achieve similar or higher performance with a similar or smaller search space size. That is, the propsed method improves the accuracy to search space size trade-off. In the case of GIN in LABEL:fig:search_space_acc_gnn, the SLT found in a 2×2$\times$2 × wider frozen source even achieves higher accuracy than those found in any dense source network.
As shown in LABEL:fig:search_space_acc_conv6 and LABEL:fig:search_space_acc_gnn, in the scenarios that only allow for medium to low pruning, there is a limit to the number of parameters that can be pruned for a given SLT sparsity, and only aggressive pruning leads to less accurate SLTs. By additionally locking parameters, our method reduces the search space further and finds SLTs of higher accuracy than those within sparse or dense networks for a similar search space size. Pruning only improves the trade-off in the scenario that allows higher pruning than locking, as shown in LABEL:fig:search_space_acc_resnet18, while our method improves the accuracy to search space trade-off in all scenarios.
@@ 4.4 Accuracy to Model Memory Size Trade-Off
The freezing mask compression scheme proposed in Section 3.2 allows to reduce the memory model size during inference by regenerating both the random weights and the freezing mask with a random number generator (RNG). Here we consider this compression and investigate the accuracy to model size trade-off offered by the SLTs found by the proposed method. Furthermore, considering the practical application of partial SLT search on specialized hardware, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization (Ramanujan et al., 2020), which can be exploited for reduced computational cost in neural engines (Hirose et al., 2022). For SKC we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al. (2020).
@@
As illustrated in Figure 7, compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy to model size trade-off in all scenarios. In other words, partial SLT search not only reduces the search space during training, but also reduces the model size during inference, further improving performance over conventional SLT search methods.
@@
Empirically, it is known that SLTs within a SKC initialized dense network achieve better performance than with continuous random weights (Ramanujan et al., 2020; Okoshi et al., 2022; López García-Arias et al., 2023; Yan et al., 2023). Our results show that such a trend can also be observed in frozen source networks. Nonetheless, in all source networks, we find that this improvement is smaller the wider the source network is, suggesting that the requirement for source networks of larger width is weaker in the case of binary weights.
Conv6 & CIFAR-1010{10}10
Method
(Source Network)
Weight
Initialization
Sparsity
Pruning
Ratio [%]
Locking
Ratio [%]
Top-1 Test
Acc. [%]
Search Space
[# Parameters]
Model Size
[MB]
Weight Training KU - - - $87.1$ $2.26$ M $8.63$
SLT (Dense) SKC $50$ 00 00 $86.2$ $2.26$ M $0.27$
SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $1.24$ M $0.15$
SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $1.13$ M $0.13$
ResNet-${18}$ CIFAR-1010{10}10
Weight Training KU - - - $92.5$ $11.16$ M $42.63$
SLT (Dense) SKC $90$ 00 00 $91.0$ $11.16$ M $1.37$
SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $1.12$ M $0.17$
GIN & OGBN-Arxiv
Weight Training KU - - - $70.1$ $0.38$ M $1.452$
SLT (Dense) SKC $20$ 00 00 $70.1$ $0.38$ M $0.054$
SLT (Sparse) SKC $20$ $15$ 00 $63.1$ $0.32$ M $0.047$
SLT (Frozen) SKC $20$ 00 $40$ $70.6$ $0.23$ M $0.036$
4.5 Result Analysis
[…]
Interestingly, the SLT found in the frozen GIN model achieves higher accuracy than the trained-weight network, even though the model size and search space are reduced by 40.3×40.3$\times$40.3 × and 1.7×1.7$\times$1.7 ×, respectively. The SLT found within a dense network also achieves comparable accuracy to the trained-weight network, but the SLT found by our method performs better and is smaller. Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem (Li et al., 2018). Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem (Huang et al., 2022) and achieve higher accuracy. As our method is even more accurate than dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
@@ 5 Conclusion
This paper capitalizes on the fact that pruning a randomly weighted network at initialization reduces the SLT search space, but identifies that it limits the search to a suboptimal sparsity region. This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search. Freezing allows to search for SLTs efficiently in the optimal sparsity range, while further reducing search space during training and model size during inference. Experimental results show that SLTs found in frozen networks improve the accuracy to search space and accuracy to model size trade-offs compared to SLTs found in dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse source networks (Gadhikar et al., 2023). Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression (Zhou et al., 2019; Wimmer et al., 2020), we identify a scenario in SLT training where it can be used for raising accuracy, even outperforming weight-trained networks. Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, and as a result, it offers an opportunity for reducing off-chip memory access in specialized SLT accelerators (Hirose et al., 2022). We hope that this work will serve as a base for raising further the efficiency of SLT training and inference.
Impact Statements
@@ Appendix A Proof for Strong Lottery Tickets (SLTs) Existence in Frozen Networks
This section describes the proofs of the lemma and theorem introduced in the manuscript. For reference, we present a theorem presented by Gadhikar et al. (2023) in advance.
Theorem A.1 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
@@
We first extend the lemma of the subset-sum approximation shown by Lueker (1998) to the case where some random variables are always included in the subset-sum.
@@
To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker (1998):
@@
From the proof of Corollary 3.1 in Lueker (1998), for any $\gamma\in(0,\frac{1}{4})$, we know that
@@
By similar arguments as in Lueker (1998) or Pensia et al. (2020), we can easily generalize Lemma A.2 to the case where the distribution followed by $X_{i}$ contains the uniform distribution. Also, by combining our proof and the proof of Lemma 2.4 in Gadhikar et al. (2023), we can prove the following extension of Lemma A.2:
@@
By replacing Lemma 2.4 in Gadhikar et al. (2023) with our Lemma A.3, we obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al. (2023). ∎

### v2

@@ 3.3 Optimal Positioning of the Search Space
Figure 5: Different prune:lock proportions of a $80\%$ freezing ratio using a Conv6.
Here we perform a preliminary investigation of the optimal positioning of the search space bounds for each given desired SLT sparsity by exploring different pruning:locking proportions with a fixed freezing ratio of the network. Figure 5 explores different configurations of a $80\%$ freezing ratio—a search space size of $20\%$ of the original dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10. As expected, the best-performing SLT of each prune:lock configuration is found at the center of the search space, which maximizes the number of candidate subnetworks. In other words, within a search space size $S$, the optimal SLT sparsity is $k=P_{r}+S/2$, where $P_{r}=\frac{\sum_{l}p^{(l)}_{p}n^{(l)}}{\sum_{k}n^{(k)}}$ is the global pruning ratio of the network. Conversely, for a given freezing ratio $F_{r}$ of the network and a desired SLT sparsity $k$, the optimal search space position is set by the pruning ratio $P_{r}=k-(1-F_{r})/2$ and the corresponding locking ratio $L_{r}=F_{r}-P_{r}$. In the cases where this would result in $P_{r}\!<\!0$ or $L_{r}\!<\!0$, we choose a best-effort approach that keeps the desired freezing ratio and sets the search space bounds to $P_{r}\!=\!0$ or $L_{r}\!=\!0$, respectively.
@@ 4 Experiments
This section demonstrates that freezing reduces the search space in a broad range of situations by evaluating it on image classification and graph node classification. We evaluate the impact of the freezing ratio on various network architectures, identifying three scenarios. Then, we explore trade-offs between accuracy, search space size, and model memory size for different network widths.
@@ 4.1 Experimental Settings
The proposed partial SLT search is evaluated on image classification using the CIFAR-1010{10}10 [19] and ImageNet [31] datasets, and on node classification using the OGBN-Arxiv [16] dataset. CIFAR-1010{10}10 and ImageNet train data are split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split. We test the models with the best validation accuracy and report the mean of three experiment repetitions for CIFAR-1010{10}10 and OGBN-Arxiv, and the result of one experiment for ImageNet. Standard deviation of experiments conducted more than once is plotted as error bars.
For image classification we employ the VGG-like Conv6 [32, 30], ResNet [13], and Wide ResNet [40] architectures, and for graph node classification the 4-layer modified GIN [37] architecture in Huang et al. [17], all implemented with no learned biases. ResNet and GIN use non-affine Batch Normalization [18], while Conv6 has no normalization layers. Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution [12].
SLTs are searched using an extension of Edge-Popup [30] that enforces the desired sparsity globally instead of per-layer [33]. On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively. On ImageNet, scores are optimized by the same setting as ResNet-1818{18}18 on CIFAR-1010{10}10, but a $256$ batch size. On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW [23] with weight decay $0.0001$ and initial learning rate $0.01$. All experiments use cosine learning rate decay [22]. These can be adequately verified with two NVIDIA H100 SXM5 94GB GPUs.
@@ 4.2 Reducing the Search Space with Freezing
Applying the much larger ResNet-1818{18}18 to the same task results in much stronger overparametrization. Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting of much higher pruning than locking. Even though $90\%$ of the search space is reduced in this scenario, we can find $90\%$ sparse SLTs with $88.1\%$ accuracy.
As an example of the scenario where optimal SLT sparsities are found in the denser range, benefiting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin. Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the search space our method finds similarly performing SLTs of $69.8\%$ accuracy with $40\%$ sparsity.
Interestingly, with low SLT sparsity (e.g., $10\%$ sparsity) in all three scenarios, despite the reduced search space, SLTs within frozen networks achieve higher accuracy than SLTs within dense networks. These results suggest that parameter freezing at an appropriate ratio has the effect of avoiding the inclusion of low-grade local optimal solutions in the search space. It is known that Edge-Popup struggles to find the SLTs of the global optimal solution [9]. While searching for SLTs in a dense network leads to convergence to a local optimal solution, a moderate reduction of the search space may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space. In other words, we conjecture that if the network is properly frozen, the local optimal solution for the reduced search space is close to the global optimal solution of the entire search space.
4.3 Accuracy-to-Search Space Trade-Off
@@
The results show in all cases that the found SLTs in wide frozen networks for a given sparsity achieve similar or higher performance with a similar or smaller search space size compared to the dense source network. That is, the proposed method improves the accuracy-to-search space size trade-off.
As shown in LABEL:fig:search_space_acc_conv6 and LABEL:fig:search_space_acc_gnn, in the scenarios that only allow for medium to low pruning, there is a limit to the number of parameters that can be pruned for a given SLT sparsity, and only aggressive pruning leads to less accurate SLTs. By additionally locking parameters, our method reduces the search space further and finds SLTs of higher accuracy than those within sparse or dense networks for a similar search space size. Pruning only improves the trade-off in the scenario that allows higher pruning than locking, as shown in LABEL:fig:search_space_acc_resnet18, while our method improves the accuracy-to-search space trade-off in all scenarios.
Figure 7: Freezing achieves better accuracy-to-search space trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks. SLT sparsity set according to highest accuracy in Figure 6 with dense source networks.
Figure 8: Compared to sparse ($\bullet$) or dense ($\bullet$) source networks (top-left is better), freezing achieves better accuracy-to-memory size trade-off.
@@ 4.4 Accuracy to Model Memory Size Trade-Off
The freezing mask compression scheme proposed in Section 3.2 allows to reduce the memory model size during inference by regenerating both the random weights and the freezing mask with a random number generator (RNG). Here we consider this compression and investigate the accuracy-to-model size trade-off offered by the SLTs found by the proposed method. Furthermore, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization [30], which can be exploited for reduced computational cost in neural engines [14]. For SKC we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al. [30].
@@
As illustrated in Figure 8, compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy-to-model size trade-off in all scenarios. In other words, partial SLT search not only reduces the search space during training, but also reduces the model size during inference, further improving performance over conventional SLT search methods.
@@
Empirically, it is known that SLTs within a SKC initialized dense network achieve better performance than with continuous random weights [30, 26, 21, 38]. Our results show that such a trend can also be observed in frozen source networks. Nonetheless, in all source networks, we find that this improvement is smaller the wider the source network is, suggesting that the requirement for source networks of larger width is weaker in the case of binary weights.
4.5 ImageNet Experiments
Figure 9: Freezing enhances the accuracy-to-model size trade-off even on a practical setting with large-scale dataset and models.
Finally, we evaluate our method using larger models on a large-scale dataset: deeper and wider ResNets on ImageNet. Since SLTs with $80\%$ sparsity achieve the highest accuracy in a dense source ResNet-5050{50}50 (for details, see Figure 10 in Appendix B), we compare the three methods using $80\%$ SLT sparsity. Figures 2 and 9 compare the accuracy-to-search space and accuracy-to-model size trade-offs, respectively, in finding the $80\%$ sparsity SLTs between the proposed and conventional methods with SKC using ImageNet. In both of the figures, such SLTs searched from non-frozen (blue), $70\%$ pruned (orange), and $65\%$ pruned and $5\%$ locked ($70\%$ frozen, green) are compared. Despite the more challenging setting and the significant $70\%$ search space reduction, our method finds SLTs that are substantially more accurate than other methods for the same search space size (Figure 2) and for the same model memory size (Figure 9). This result demonstrates that the effective combination of parameter pruning and locking at initialization can improve the SLT search efficiency even on large-scale datasets and models. Notably, Figure 9 reveals that the proposed method enhances the accuracy-to-model size trade-off, validating our claim it could be deployed on specialized inference hardware, such as the one like Hirose et al. [14], to enhance the efficiency of practical SLT applications.
5 Limitations
Our results have the following limitations: 1) Although the proposed method can effectively reduce the search space for training and the memory model size for inference in principle, actual hardware implementation remains for future work. The previous work of the SLT-specialized neural engine, Hiddenite [14], would be a promising direction for such future implementation. 2) This paper provides theoretical support for the existence of SLTs in frozen networks, but it does not provide their superiority in the approximation capability compared to random pruning, a special case of random freezing, which may depend on the sparsity nature of target networks. Also, the theoretical consideration of the appropriate proportion of these ratios is left for future work. 3) Even though we demonstrate our method on various model architectures including CNNs, ResNet families and GIN, following previous SLT work [41, 30, 21, 17, 38], we have to leave it for future work to apply our method to Transformers [35] because SLT itself has not yet been established for Transformers.
@@ 6 Conclusion
This paper capitalizes on the fact that pruning a randomly weighted network at initialization reduces the SLT search space, but identifies that doing so limits the search to a suboptimal sparsity region. This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search. Freezing allows to search for SLTs efficiently in the optimal sparsity range, while further reducing the search space during training and the model size during inference. Experimental results show that SLTs found in frozen networks improve the accuracy-to-search space and accuracy-to-model size trade-offs compared to SLTs found in dense [30, 33] or sparse source networks [10]. Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression [41, 36], we identify scenarios in SLT training where they can be used for raising accuracy. Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, offering an opportunity for reducing off-chip memory access in specialized SLT accelerators [14]. We hope this work will serve as a base for raising the efficiency of SLT training and inference.
Acknowledgments and Disclosure of Funding
This work was supported in part by JSPS KAKENHI Grant Number JP23H05489.
@@ Appendix A Proof for Strong Lottery Tickets (SLTs) Existence in Frozen Networks
This section describes the proofs of the lemma and theorem introduced in the manuscript. For reference, we present a theorem presented by Gadhikar et al. [10] in advance.
Theorem A.1 (SLT Existence in Sparse Networks [10]).
@@
We first extend the lemma of the subset-sum approximation shown by Lueker [24] to the case where some random variables are always included in the subset-sum.
@@
To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker [24]:
@@
From the proof of Corollary 3.1 in Lueker [24], for any $\gamma\in(0,\frac{1}{4})$, we know that
@@
By similar arguments as in Lueker [24] or Pensia et al. [28], we can easily generalize Lemma A.2 to the case where the distribution followed by $X_{i}$ contains the uniform distribution. Also, by combining our proof and the proof of Lemma 2.4 in Gadhikar et al. [10], we can prove the following extension of Lemma A.2:
@@
By replacing Lemma 2.4 in Gadhikar et al. [10] with our Lemma A.3, we obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al. [10]. ∎
Appendix B Additional Experimental Results
B.1 Accuracy of SLT within Dense ResNet-50
This section introduces the preliminary experiment used for determining the main experimental setup. Figure 10 compares the accuracy of SLTs within a dense ResNet-5050{50}50 source at different sparsity using ImageNet. The SLT with $80\%$ sparsity is the most accurate, reaching $66.8\%$ accuracy.
Figure 10: Accuracy comparison of SLTs of different sparsity within a dense ResNet-5050{50}50 on ImageNet.
B.2 Result Analysis
[…]
Interestingly, the SLT found in the frozen GIN model achieves comparable accuracy to the trained-weight network, even though the model size and search space are reduced by 40.3×40.3$\times$40.3 × and 1.7×1.7$\times$1.7 ×, respectively. The SLT found within a dense network also achieves comparable accuracy to the trained-weight network, but the SLT found by our method is smaller. Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem [20]. Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem [17] and achieve higher accuracy. As our method is similarly accurate compared to dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
Although the ImageNet experiment falls into the scenario where a relatively sparse SLT is more accurate (like the ResNet-1818{18}18 experiment on CIFAR-1010{10}10), the SLT found in a sparse network is significantly less accurate. On the other hand, the inclusion of parameter locking shows a significant improvement in SLT accuracy. These results suggest that the number of highly accurate SLT patterns decreases as the difficulty of the problem increases, and the way the strategy of search space reduction affects the performance of the found SLTs more severely. As also shown in Figure 2 in the main manuscript, the SLT found in the frozen Wide ResNet-5050{50}50 achieves $1.4$ points higher accuracy than the SLT found in the dense ResNet-3434{34}34, even though the search space is smaller, indicating superiority of our method in improving the accuracy-to-search space trade-off.
Conv6 & CIFAR-1010{10}10
Method
(Source Net.)
Weight
Init.
Sparsity
Pruning
Ratio [%]
Locking
Ratio [%]
Top-1 Test
Acc. [%]
Search Space
[# Param.]
Model
Size [MB]
Weight Training KU - - - $87.1$ $2.26$ M $8.63$
SLT (Dense) SKC $50$ 00 00 $86.2$ $2.26$ M $0.27$
SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $1.24$ M $0.15$
SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $1.13$ M $0.13$
ResNet-${18}$ CIFAR-1010{10}10
Weight Training KU - - - $92.4$ $11.16$ M $42.63$
SLT (Dense) SKC $90$ 00 00 $91.0$ $11.16$ M $1.37$
SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $1.12$ M $0.17$
GIN & OGBN-Arxiv
Weight Training KU - - - $70.1$ $0.38$ M $1.452$
SLT (Dense) SKC $20$ 00 00 $70.0$ $0.38$ M $0.054$
SLT (Sparse) SKC $20$ $15$ 00 $62.3$ $0.32$ M $0.047$
SLT (Frozen) SKC $20$ 00 $40$ $69.2$ $0.23$ M $0.036$
ResNet-${50}$ ImageNet
Weight Training KU - - - $74.4$ $25.50$ M $97.49$
SLT (Dense) SKC $80$ 00 00 $66.8$ $25.50$ M $3.24$
SLT (Sparse) SKC $80$ $70$ 00 $47.1$ $7.65$ M $1.11$
SLT (Frozen) SKC $80$ $65$ $5$ $55.2$ $7.65$ M $1.11$
ResNet-${34}$ ImageNet
SLT (Dense) SKC $80$ 00 00 $62.5$ $21.78$ M $2.66$
SLT (Sparse) SKC $80$ $70$ 00 $40.7$ $6.53$ M $0.84$
SLT (Frozen) SKC $80$ $65$ $5$ $51.2$ $6.53$ M $0.84$
ResNet-${18}$ ImageNet
SLT (Dense) SKC $80$ 00 00 $54.0$ $11.68$ M $1.43$
SLT (Sparse) SKC $80$ $70$ 00 $25.2$ $3.50$ M $0.45$
SLT (Frozen) SKC $80$ $65$ $5$ $39.4$ $3.50$ M $0.45$
Wide ResNet-${50}$ ImageNet
SLT (Dense) SKC $80$ 00 00 $70.8$ $68.81$ M $8.46$
SLT (Sparse) SKC $80$ $70$ 00 $57.7$ $20.64$ M $2.72$
SLT (Frozen) SKC $80$ $65$ $5$ $63.9$ $20.64$ M $2.72$
Appendix C Impact Statements